OpenStudy (anonymous):

log₅(x-3)+log₅x=log₅4 how do i solve this?

OpenStudy (shubhamsrg):

base is 5 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

u know the properties of logarithm, @mathnoob1 ?

OpenStudy (shubhamsrg):

well doesnt matter.. use this,,log a +log b = log ab and if log m = log n then m=n try now note : the bases should be equal here..

OpenStudy (anonymous):

if apply the properties, u will get x(x-3)=4 now, solve this

OpenStudy (anonymous):

x²-3x=4 ???

OpenStudy (anonymous):

factor your equation

OpenStudy (anonymous):

roots will be 4 and -1 but u have to discard -1 as log of negative number is not defined so, the answer will be 4

OpenStudy (anonymous):

hmm icic thank you everyone =)

OpenStudy (anonymous):

@ninhi5 , u r wrong

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

how do u define log(-2), @ninhi5 ?

OpenStudy (anonymous):

@Arnab09 is right.

OpenStudy (anonymous):

x.(x-3)=4

OpenStudy (anonymous):

well guy the answer for this is 4.. and aranb09 answer it right

OpenStudy (anonymous):

x>0 and x-3>0

OpenStudy (anonymous):

checking doesn't matter

OpenStudy (anonymous):

so x=4

OpenStudy (anonymous):

guy this kind of equation can i actually use quadratic equation to solve it?

OpenStudy (anonymous):

haha it's late at night

OpenStudy (anonymous):

gr8.. good night